OpenStudy (anonymous):

Divide 35b^5 + 20ab^3 + 20a^2b^2 by 5b^2. What is the quotient? Select one of the options below as your answer: A. 7b^3 - 4ab - 4a^2 B. 7b^3 + 4ab + 4a^2 C. 7b^7 + 4ab^5 + 4a^2b^4 D. 35b^3 + 20ab + 20a^2

OpenStudy (lalaly):

B

OpenStudy (anonymous):

b

OpenStudy (anonymous):

35b^5/5b^2=7b^3; 20ab^3/5b^2=4ab; 20a^2b^2/5b^2=4a^2 OR 7b^3+4ab+4a^2; answer B

OpenStudy (anonymous):

It's b